OpenStudy (anonymous):

Explain, in complete sentences, how you would divide the following expression: (2 points) index10 sqrt x^7 / index7 sqrt x^4 Part 2: Write the simplified answer in radical form. (

OpenStudy (anonymous):

Apply common denominator: 7/ 10 and 4/7

OpenStudy (anonymous):

= index 40 root x^9

jhonyy9 (jhonyy9):

so the easy way i think that is if you write the radical terms in form of exponents - than will get index10 sqrt(x^7) = x^(7/10) - so because you know that the denominator of fraction of exponent sign the index of radical allways - so than index 7sqrtx^4 = x^(4/7) - and hence from your exercise you will get x^(7/10) ---------- - so what you can write it - because you know that when there are two the x^(4/7) same terms like in our case x with exponents in fraction so than you need just subtract the exponents for you can getting the result ,so like : x^(7/10) ---------- = x^(7/10 - 4/7) = x^((49-40)/70) = x^(9/70) x^(4/7)

jhonyy9 (jhonyy9):

- so hope so much that is understandably sure right good luck bye